could you please help me in Solving using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.

-7x + 5y = -56
-21x – 2y = -168
greatly appreciated

Answer by lwsshak3(11628)

could you please help me in Solving using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.
..
-7x + 5y = -56
-21x – 2y = -168
..
multiply first equation by 3
-21x + 15y = -168
-21x – 2y = -168
subtract
17y=0
y=0
-7x+5y=-56
7x+0=56
x=56/7
x=8
solution: (8,0)